OpenStudy (anonymous):

what other expression is equivalent to (7x-5) - (3x-2)?

OpenStudy (agent47):

just simplify it: \[(7x-5)-(3x-2)=7x-5-3x+2=7x-3x-5+2=4x-3\]

OpenStudy (anonymous):

Thank you